OpenStudy (anonymous):

Is the equation 3+z=-19 true, false or open?

jimthompson5910 (jim_thompson5910):

Any ideas?

OpenStudy (anonymous):

No i'm stuck

jimthompson5910 (jim_thompson5910):

hint: when they say "is the equation true" they mean "Is it ALWAYS true for ALL values of z?"

jimthompson5910 (jim_thompson5910):

So is 3+z=-19 always true for all values of z?

OpenStudy (anonymous):

no?

jimthompson5910 (jim_thompson5910):

good, there's only one value of z that works, do you know what it is?

OpenStudy (tkhunny):

This is known as "Conditionally True".

jimthompson5910 (jim_thompson5910):

subtract 3 from both sides to get 3+z=-19 3+z-3=-19-3 z = -22 So it's only true when z = -22. This means that it's neither 100% true nor 100% false. So it's an open statement or sentence.